# Win Ratio with Multiple Thresholds for Composite Endpoints

### v4

@@ Acknowledgement
Appendix
Appendix A Generalization of WR-MT with additional thresholds and multiple non-fatal events
In the main text, we introduce the 4-stage WR-MT with survival time and time-to-hospitalization endpoints. For the ease of generalization, we first rewrite each stage in $\text{WR-MT}(d,t_{1},0,0)$ with comparison functions:
@@ Appendix A Generalization of WR-MT with additional thresholds and multiple non-fatal events
The WR-MT formed by the survival time (with one non-zero threshold $d$) and $K$ nonfatal events can be decomposed into the following $2(K+1)$ components:
@@
Based on the above two scenarios, in the most general scenario, for survival time and $K\geq 1$ nonfatal events, we may specify $L\geq 1$ non-zero thresholds for each of them to form the $(K+1)(L+1)$-stage WR-MT. Denote non-zero thresholds for survival time as $\left\{d_{1},d_{2},...,d_{L}>0\right\}$, and for the $k$th nonfatal event as $\left\{t_{k1},t_{k2},...,t_{kL}>0\right\}$. The general expression for the win-loss score of pair comparison between $i$ and $j$ becomes:
@@
Appendix B Additional simulation scenarios
In this appendix, we include simulation results for 8 additional simulation scenarios S9-S16, as listed in Table A-1. The simulation setup is the same as that in the main text for simulation scenarios S1-S8. The table version of the power achieved by WR-AT and WR in simulation scenarios S1-S8 is also included.
Table A-1: Settings of treatment effect sizes and correlations between survival time and time-to-hospitalization endpoints in eight additional simulation scenarios
@@ Appendix B Additional simulation scenarios
B.1 Performance of WR-AT
In simulation scenarios S9&S10, where the weak treatment effect is on the survival time endpoint only, compared to WR, WR-AT achieves lower but comparable power when correlation is not assumed, and the difference in power is larger with correlation assumed. Such observation is similar to that in simulation scenarios S3&S4. In simulation scenarios S11-S14, WR-AT and WR perform well as the treatment effects are on both endpoints and the overall strength of such effects is strong enough. In simulation scenarios S15&S16, where very weak treatment effects are on both endpoints, WR-AT achieves comparable but slightly higher power than WR. This is likely a result of providing more chances for the time-to-hospitalization endpoint, where the treatment effect is easier to observe, as the administrative censoring rate is usually lower than that for the survival time endpoint.
Table A-2: Empirical power (%) obtained by WR-AT and WR in simulation scenarios S1-S16
Scenario FU WR WR-AT Scenario FU WR WR-AT
@@ B.1 Performance of WR-AT
B.2 Influence of caliper
In simulation scenarios S9-S16, the performance of WR-AT is generally comparable under different caliper choices, including the combined caliper strategy, and those differences observed are as expected. In simulation scenario S9, increased caliper values and adopting the combined caliper strategy achieve lower power for addressing the time-to-hospitalization endpoint with spurious negative “treatment effects”. In simulation scenarios S15&S16, on the contrary, higher power is achieved by them due to the easier-to-observe treatment effects on the time-to-hospitalization endpoint.
Table A-3: Empirical power (%) obtained by WR-AT with different calipers $c$ in simulation scenarios S9-S16
@@ B.2 Influence of caliper
B.3 Influence of weight
In simulation scenarios S9-S16, the performance of WR-AT is generally comparable under different weight choices, including the combined caliper strategy, and those differences observed are as expected. In simulation scenario S9, increased weights achieve lower power for addressing the time-to-hospitalization endpoint with spurious negative “treatment effects”. In simulation scenarios S15&S16, on the contrary, higher power is achieved by increasing the weight due to the easier-to-observe treatment effects on the time-to-hospitalization endpoint.
Table A-4: Empirical power (%) obtained by WR-AT with different weights $w$ in simulation scenarios S9-S16
Scenario FU WR-AT(0.1) WR-AT(0.3) WR-AT(0.5) WR-AT(1.0)
@@ B.3 Influence of weight
Appendix C Influence of correlation
In this appendix, we provide additional analysis of the influence that the correlation between survival time and time-to-hospitalization endpoints has on the power. The overall win ratios are decomposed to endpoint-level, i.e., win ratios for the survival time and time-to-hospitalization endpoints. In WR, comparison results in stages 1&2 are attributed to survival time and time-to-hospitalization endpoints correspondingly. In WR-AT, stages 1&3 are attributed to the survival time endpoint, and stages 2&4 are attributed to the time-to-hospitalization endpoint. In simulation scenario S3, both WR-AT and WR have their time-to-hospitalization endpoint win ratios less than 1, as opposed to those win ratios close to 1 in simulation scenario S4. Such comparison indicates the existence of spurious negative “treatment effects” introduced by the marginal correlation between survival time and time-to-hospitalization endpoints.
Table A-5: Endpoint-level decomposition of win ratios in simulation scenarios S3&S4 averaged across 2000 replicates
Scenario FU WR WR-AT
@@ Appendix C Influence of correlation
Appendix D Additional details of the case study
The number of patients within each stratum in the case study is presented in Table A-6.
Table A-6: Number of patients within each stratum

### v5

@@ Acknowledgement
Supporting Information
Appendix S1 Generalization of the FS-MT test with additional thresholds and multiple non-fatal events
In the main text, we introduce a 4-stage FS-MT test with survival time and time-to-hospitalization endpoints. For the ease of generalization, we first rewrite each stage in $\text{FS-MT}(d,t_{1},0,0)$ with comparison functions:
@@
The FS-MT test formed by the survival time (with one non-zero threshold $d$) and $K$ nonfatal events can be decomposed into the following $2(K+1)$ components:
@@
Based on the above two scenarios, in the most general scenario, for survival time and $K\geq 1$ nonfatal events, we may specify $L\geq 1$ non-zero thresholds for each of them to form the $(K+1)(L+1)$-stage FS-MT test. Denote non-zero thresholds for survival time as $\left\{d_{1},d_{2},...,d_{L}>0\right\}$, and for the $k$th nonfatal event as $\left\{t_{k1},t_{k2},...,t_{kL}>0\right\}$. The general expression for the win-loss score of pair comparison between $i$ and $j$ becomes:
@@
Appendix S2 Additional simulation scenarios
In this appendix, we include simulation results for 8 additional simulation scenarios S9-S16, as listed in Table S1. The simulation setup is the same as that in the main text for simulation scenarios S1-S8. The table version of the power achieved by FS-AT and FS tests in simulation scenarios S1-S8 is also included.
Table S1: Settings of treatment effect sizes and correlations between survival time and time-to-hospitalization endpoints in eight additional simulation scenarios
@@
S2.1 Performance of FS-AT
In simulation scenarios S9&S10, where the weak treatment effect is on the survival time endpoint only, compared to the FS test, the FS-AT test achieves lower but comparable power when correlation is not assumed, and the difference in power is larger with correlation assumed. Such observation is similar to that in simulation scenarios S3&S4. In simulation scenarios S11-S14, FS-AT and FS tests perform well as the treatment effects are on both endpoints and the overall strength of such effects is strong enough. In simulation scenarios S15&S16, where very weak treatment effects are on both endpoints, the FS-AT test achieves comparable but slightly higher power than the FS test. This is likely a result of providing more chances for the time-to-hospitalization endpoint, where the treatment effect is easier to observe, as the administrative censoring rate is usually lower than that for the survival time endpoint.
Table S2: Empirical power (%) obtained by FS-AT and FS tests in simulation scenarios S1-S16
Scenario FU FS FS-AT Scenario FU FS FS-AT
@@
S2.2 Influence of caliper
In simulation scenarios S9-S16, the performance of the FS-AT test is generally comparable under different caliper choices, including the combined caliper strategy, and those differences observed are as expected. In simulation scenario S9, increased caliper values and adopting the combined caliper strategy achieve lower power for addressing the time-to-hospitalization endpoint with spurious negative “treatment effects”. In simulation scenarios S15&S16, on the contrary, higher power is achieved by them due to the easier-to-observe treatment effects on the time-to-hospitalization endpoint.
Table S3: Empirical power (%) obtained by the FS-AT test with different calipers $c$ in simulation scenarios S9-S16
@@
S2.3 Influence of weight
In simulation scenarios S9-S16, the performance of the FS-AT test is generally comparable under different weight choices, including the combined caliper strategy, and those differences observed are as expected. In simulation scenario S9, increased weights achieve lower power for addressing the time-to-hospitalization endpoint with spurious negative “treatment effects”. In simulation scenarios S15&S16, on the contrary, higher power is achieved by increasing the weight due to the easier-to-observe treatment effects on the time-to-hospitalization endpoint.
Table S4: Empirical power (%) obtained by the FS-AT test with different weights $w$ in simulation scenarios S9-S16
Scenario FU FS-AT(0.1) FS-AT(0.3) FS-AT(0.5) FS-AT(1.0)
@@
Appendix S3 Influence of correlation
In this appendix, we provide additional analysis of the influence that the correlation between survival time and time-to-hospitalization endpoints has on the power. The overall NB and WR are decomposed to endpoint-level, i.e., NB and WR for the survival time and time-to-hospitalization endpoints. Because the merging of multiple stages corresponding to one endpoint is non-trivial for WO due to the specific handling of ties, WO is not reported at the endpoint level. In the FS test, comparison results in stages 1&2 are attributed to survival time and time-to-hospitalization endpoints correspondingly. In the FS-AT test, stages 1&3 are attributed to the survival time endpoint, and stages 2&4 are attributed to the time-to-hospitalization endpoint. In simulation scenario S3, both FS-AT and FS tests have their time-to-hospitalization endpoint NB being negative, as opposed to those NB close to 0 in simulation scenario S4. Similarly, in simulation scenario S3, both FS-AT and FS tests have their time-to-hospitalization endpoint WR less then 1, as opposed to those WR close to 1 in simulation scenario S4. Such a comparison indicates the existence of spurious negative “treatment effects” introduced by the correlation between survival time and time-to-hospitalization endpoints.
Table S5: Endpoint-level decomposition of NB(%) in simulation scenarios S3&S4 averaged across 2000 replicates
Scenario FU FS FS-AT
S3 250 2.35 4.42 -2.07 2.00 4.02 -2.02
S3 500 5.15 7.55 -2.40 4.24 6.89 -2.65
S3 750 7.58 9.72 -2.14 6.19 9.00 -2.81
S3 1000 9.52 11.26 -1.74 7.78 10.58 -2.80
S3 1250 10.98 12.34 -1.36 9.04 11.76 -2.72
S3 1500 12.07 13.10 -1.04 10.00 12.65 -2.65
S4 250 4.26 4.29 -0.03 3.88 3.91 -0.03
S4 500 7.37 7.41 -0.04 6.59 6.61 -0.02
S4 750 9.58 9.62 -0.04 8.65 8.69 -0.04
S4 1000 11.12 11.14 -0.03 10.23 10.25 -0.02
S4 1250 12.22 12.23 -0.01 11.40 11.42 -0.02
S4 1500 13.01 13.01 -0.01 12.30 12.31 -0.01
Table S6: Endpoint-level decomposition of WR in simulation scenarios S3&S4 averaged across 2000 replicates
Scenario FU FS FS-AT
@@
Appendix S4 Additional details of the case study
The number of patients within each stratum in the case study is presented in Table S7.
Table S7: Number of patients within each stratum